OpenStudy (anonymous):

Help please!! find angle x in: 3 cot x- 2 cos=0

OpenStudy (thomas5267):

I thought you did this question already...

OpenStudy (anonymous):

I'm not able to do it.. It's quite confusing..

OpenStudy (thomas5267):

\[ \begin{align*} 3\cot(x)-2\cos(x)&=0\\ \frac{3\cos(x)}{\sin(x)}-2\cos(x)&=0\\ \cos(x)\left(\frac{3}{\sin(x)}-2\right)&=0 \end{align*} \] Try finish it.

OpenStudy (anonymous):

I think that I will be able to complete it now Thank you for helping me.. :)